OpenStudy (anonymous):

integral of (ln^3(2x)) / (x)

OpenStudy (anonymous):

For clarification \[\int\limits_{?}^{?}\frac{ \ln ^{3}2x }{ x}\]

OpenStudy (anonymous):

did u try u-substitution

OpenStudy (anonymous):

yeah, i got \[\frac{ \ln^{4}2x }{ 4 }+ c\]

OpenStudy (anonymous):

Is it right? :O

OpenStudy (raden):

that's right... the processing integral use 2 times the u-sub right ?

OpenStudy (anonymous):

umm i just used it once then manipulated the integral to make (1/x)(dx) into (2dx/2x) which is du

OpenStudy (raden):

nope, the result is same ^^

OpenStudy (anonymous):

thanx